# Extraction of Medication and Temporal Relation from Clinical Text by Harnessing Different Deep Learning Models

### v1

@@ Abstract
1 Introduction
With the continuous growth of the need for medical text mining, extraction of temporal information has become a significant field of Natural Language Process (NLP) Su et al. (2004), e.g. when the disease was diagnosed, how long a patient should take some medicines. Due to its importance for many other NLP tasks, temporal information tagging and annotation from clinical texts have become an active research field (Figure 1 for example). Clinical texts including electronic health records (EHRs) consist of medical history, diagnoses, medications, treatment plans, immunisation dates, allergies, radiology images, and laboratory and test results, which contain a lot of useful information that can provide the opportunity to support medical research Han et al. (2023). These include the study of the side effects of medications (a.k.a adverse drug effects) taken over time, helping doctors make better decisions for patients, and finding the correlation between drug use and its outcomes, etc. Fredriksen et al. (2014); Alrdahi et al. (2023).
@@ 1 Introduction
There are two major sub-tasks: 1) Medication Entity Recognition, which involves finding and identifying named drugs mentioned in unstructured text, the first step toward automating the extraction of medications from EMRs. 2) Automatic Temporal Relation Extraction from free text EMRs. We approach these two steps by applying several advanced DL methods including the BiLSTM+CRF, CNN-BiLSTM, and BERT-base-CNN models to train on the available data sets. Additionally, a SparkNLP tool (DateNormalizer) Kocaman and Talby (2021) is used to extract and normalise dates from relative date-time phrases.
2 Methodological Design
Our designed methodology includes five main steps as shown in Figure 2, which are Data Pre-processing, Word Embedding, Modelling, Time Expression and Event Extraction, and Temporal Relations and Medications Candidates Classification. The Step-1 on data pre-processing includes sentence segmentation, tokenisation, Part of Speech (POS) tagging, and syntactic parsing, which details are located in Appendix with examples (A: Pre-Processing Details: “A: Pre-Processing Details”). The goal of data pre-processing is to produce structured text with features. Regarding Step-2, we tried three different kinds of word embedding techniques including one-hot encoding, bag-of-words, and GloVe embedding, which are detailed in Appendix (B: Word Embeddings: “B: Word Embeddings”).
@@ 2 Methodological Design
2.1 Deployed Learning Models
2.1.1 BiLSTM
@@
In machine learning fields, the problem that the model cannot get the previous moment information due to the vanishing gradient is also called the long-term dependence problem. In order to solve this problem, many researchers have tried to improve the Recurrent Neural Network (RNN) model, including BRNN Arisoy et al. (2015), GRU Tang et al. (2016), and LSTM Hochreiter and Schmidhuber (1997), etc., among which the LSTM model is the most widely used one.
@@
LSTM is an improved RNN structure that can effectively avoid the long-term dependence problem that the traditional RNNs cannot solve. In the later work, many people adjusted and popularised the model, and the modern, widely used LSTM structure came into being. LSTM network has the same chain structure as traditional RNNs, but the repeating module has a different design. The RNN model only has a single layer of repeated modules. In contrast, the LSTM network has four interacting neural network layers, which makes the LSTM model more complex and sacrifices a certain amount of computational time, but the model output becomes better. The standard LSTM model is shown in Figure 7 (Appendix C: More Details on Learning Models), which mainly includes three groups of adaptive element multiplication gates, namely, the input gate, forgetting gate, and output gate.
@@
In actual text processing applications, the text is characterised by sequential feature correlation, and words in the current moment are not only affected by words in the past moment but also by words in the future moment. In order to make the model contain contextual information at the same time, researchers put forward a bidirectional LSTM model (BiLSTM) Zhang et al. (2015); Zhou et al. (2016). BiLSTM contains two LSTM layers, which train the forward and backward sequences, respectively, and both LSTM layers are connected to the output layer. This forward-backward two-layer LSTM structure can provide historical information and future information to the output layer at the same time, which can not only retain the advantages of LSTM to solve long-term dependence but also take into account the context information and effectively deal with sequence problems.
2.1.2 CRF model
@@
Conditional random fields (CRFs) Lafferty et al. (2001) are a type of statistical modelling methodology frequently employed in pattern recognition and machine learning, as well as for structured prediction, which is a widely-applied modelling strategy for many NLP tasks Han et al. (2013, 2015). Data having an underlying graph structure can be better modelled using CRFs, a graphical model that can take advantage of structural dependencies between outputs. In contrast to classifiers, which make label predictions for isolated data without considering their “neighbouring” samples, CRFs are able to account for context Maldonado et al. (2017); Moreau et al. (2018); Wu et al. (2022). In order to accomplish this goal, the predictions are modelled in the form of a graphical model, which depicts the existence of relations between the predictions. The choice of the graph to implement is context-specific. In NLP fields, for instance, “linear chain” CRFs are prevalent because each prediction depends only on its immediate neighbours. Connecting neighbouring and/or analogous regions ensures that all regions within a picture are predicted in the same way by the graph. The mathematical calculation of CRFs is detailed in Appendix (C: More Details on Learning Models).
2.1.3 CNN model
@@
In deep learning (DL), a classic convolutional neural network (CNN) is a class of artificial neural networks (ANN), most commonly applied to analyse visual imagery Albawi et al. (2017). A CNN model mainly consists of these layers: input layer, convolution layer, relu layer, pooling layer and full connection layer. One of CNN’s most alluring quality is its ability to make use of spatial or temporal correlation in data. CNN is divided into multiple learning phases, each of which is made up of a combination of convolutional layers, nonlinear processing units, and subsampling layers Goodfellow et al. (2016). In a CNN, each layer of the network uses a set of convolutional kernels to perform many transformations. The convolution process simplifies extracting useful features from spatially correlated data points. More on CNNs are displayed in Appendix C.
2.1.4 BERT model
@@
Brief for “Bidirectional Encoder Representations from Transformers”, BERT Devlin et al. (2018) is founded on Transformer structure Vaswani et al. (2017), a cutting-edge DL model in which all output elements are related to all input elements and the weights between them are dynamically computed based on their connections. BERT is unique in that it can be used with text that is read either from left to right or right to left. With this characteristic, BERT was originally taught for two NLP tasks: Next Sentence Prediction and Masked Language Modelling. The input for BERT is comprised of Token Embeddings, Segment Embeddings, and Position Embeddings, followed by pre-training as indicated in the figure 11 (Appendix).
The final output of the [CLS] parameter of the BERT is set as C, the input embedding is designated as E, and the final output of the ith token is designated as T. This will enable effective dynamic feature encoding of polysemous words. In other words, the same word might produce a variety of word vector outputs depending on the language environment. Additionally, the pre-trained model’s embedding layer transfers the knowledge from the corpus to make the word vectors more generic, greatly enhancing the model’s accuracy.
Instead of using word embedding, the conventional method of language processing, BERT assigns each word to a vector that only conveys a single dimension of its meaning. Extensive quantities of labeled data are required for word embedding models. However, because all words are in some sense bound to a vector or meaning, they have difficulty with the context-heavy, predictive nature of question responding. In order to prevent the target word from “seeing itself” (i.e., developing a meaning apart from its context), BERT employs a technique called masked language modelling. The advent of pre-trained language models in recent years has marked a significant milestone in developing NLP. Pre-trained language model methods from the past focused on acquiring high-quality word embeddings. Word vectors are context-free, notwithstanding their ability to capture a word’s semantics. Since this is the case, individuals are unable to understand the deeper meanings included in the text, such as its syntax and semantics.
2.2 Time Expressions and Events Named Entity Recognition
@@
The goal of NER task is to mark up temporal information present in clinical text in order to enable reasoning on relevant events (medications) for each patient. In this work, a commonly used format in entity tagging “Inside–outside–beginning (IOB)” Ramshaw and Marcus (1999), is used, which is explained below:
@@
2.3 Temporal Relations and Medication Candidates Classification
Relation extraction is the task of predicting attributes and relations for entities in a sentence. Extracted relations usually occur between two or more entities of a particular type (e.g. Medication, Dosage) and fall into several semantic categories. In this work, these relations that are discussed are between medications and corresponding dates (e.g. “He has been treated with simvastatin since 12/04/1998”, which contains the relation “simvastatin” happened after “12/04/1998”). Generally, there are eight types of temporal relations, including before, after, simultaneous, overlap, begun_by, ended_by, during, and before_overlap suggested by Sun et al. (2012). However, only the types of before, after, and overlap meet our task. A BERT-based model with CNN as an additional learning layer is implemented for this step.
2.4 Post-processing
After the phrases of entity recognition and relation extraction, to make the result more robust and representative, this post-processing step is to convert results to a structured table or CSV shape. However, in the previous 2012 i2b2 NLP challenge on temporal relation extraction, post-processing was considered a set of classification routines for the EVENTs only, and the TIMEx catergory remained unclassified Lin et al. (2013). In this work, we apply an open resource tool SparkNLP [^1] https://github.com/JohnSnowLabs/spark-nlp, combined with our medication entity recognition and temporal relation extraction system to generate more informative results and a structured representation table e.g. the example Table 1 generated from the example text in Figure 1.
@@ 2.4 Post-processing
Table 1: Extraction Result Example
3 Experimental Evaluations
3.1 Dataset and Evaluation Metrics
In this work, the i2b2 challenge track 2009 and 2012 datasets were used. The objective is to gather details about every drug that the patient is known to take for each patient report that is given. Partners Healthcare’s discharge summaries will be the input for the medication challenge Sun et al. (2012). These files are loosely organised. Many drugs are covered in the narrative text in addition to being listed by name in lists of pharmaceuticals. According to statistical analysis, it contains 447 EMRs, with 252 annotated, which could be used to train and test the model. While the data from the i2b2-2012 temporal relations challenge has similar content as the challenge track of 2009 does, it has a different format, additional entity categories, and temporal relations between events. It consists of 760 files including 190 TXT files (texts), 190 XML.EXTENT files (annotation and position in original text), XML files (annotation and text) and TLINK files (relations) Gurulingappa et al. (2012). The data sets are divided into 70%, 15% and 15% for training, validation, and testing, respectively on both these two i2b2 challenges. The evaluation metrics we used are Precision, Recall, F1 score, and Accuracy based on exact matching, detailed in Appendix (“E: Dataset and Metrics”) together with dataset analysis.
3.2 Model-I: BiLSTM-CRF for NER Task
Combined with the two models BiLSTM and CRF, we carried out the NER task of electronic diseases using BiLSTM-CRF. The specific steps carried out are as follows:
@@ 3.2 Model-I: BiLSTM-CRF for NER Task
Using one-hot word embedding, the evaluation scores on 2009 and 2012 data sets are shown in Table 2 and 3 with the explanations of each tag in Table 9 and 10. The parameter sets for these two experiments as well as the learning curves on Accuracy and Loss changes are listed in Appendix (F: Model Analysis).
Table 2: NER Evaluation on the 2009 Dataset with BiLSTM-CRF.
@@
Table 3: NER Evaluation on the 2012 Dataset with BiLSTM-CRF.
@@
As mentioned, evaluation metrics were based on exact-matching criteria and with this setting, the BiLSTM+CRF model achieves a 0.98 weighted average accuracy and 0.69 of Macro averaging. Apparently, combined with the evaluation scores of each subcategory, it can be seen that the amount of entities of each category varies significantly due to the uneven data distribution, which leads to the scores between entity categories varying wildly.
As demonstrated in the result of 2009 data, this BiLSTM+CRF model has better prediction performance on class “o” and “B-m” in the 2009 dataset. In general, the BiLSTM+CRF model shows better performance on the 2009 dataset than the 2012 dataset for the NER task.
In order to explore the impacts of different embedding inputs for the BiLSTM+CRF model, we trained the model with Glove (glove.6B.50d), word2vector, in addition to one-hot encoding, and the average and standard deviations of metrics are recorded. Table 4 demonstrates the performance of each model on the same dataset. Glove embedding achieved 72.48% of precision, 52.58% of recall, and 60.86% of f1-score, respectively which increased precision by around 2%. For word2vec embedding, this model achieved exact 73.62% precision, 60.34% of recall, and 61.48% of f1-score achieving an further improvement of 1.14% in precision, an increase of 7.76% in recall and an improvement of 0.62% than Glove embedding.
Table 4: Macro Avg Results of Different Embeddings on the 2009 Test Set using BiLSTM+CRF.
@@
Table 5: Macro Avg Results of Experiments on the 2009 Test Dataset with CNN-BiLSTM.
@@
For the overall trend, we observed that the prediction performance really depends on the amount of sub-category data for training, which is that the more data, the more information obtained, leading to a better evaluation score. For entity categories with a smaller number of entries, the recognition effect is poorer due to insufficient model learning. For example, sub-classes B-r and I-du have lower precisions than B-m and B-do in the 2009 dataset. However, there are some exceptions that with more training data samples, the model still shows poorer results for sub-class B-r than sub-class B-du.
In addition, the results based on precision, recall, and F1-score are pretty different. For most sub-classes, the model has higher precision than recall values which means it returns fewer results, but most of its predicted labels are correct when compared to the gold labels. Namely, the percentage of Negative labels is lower. In conclusion, the BiLSTM+CRF model can not fully solve the problem of unbalanced data distribution. On the dataset with unbalanced data distribution, the performance for the entity category with a smaller number of samples is lower.
3.3 Model-II: CNN-BiLSTM for NER

### v2

# Extraction of Medication and Temporal Relation from Clinical Text using Neural Language Models

@@ Abstract
Index Terms:
Neural Language Models, Deep Learning Models, Natural Language Processing, AI for Healthcare, Medication and Temporal Relation
I Introduction
With the continuous growth of the need for medical text mining, extraction of temporal information has become a significant field of Natural Language Process (NLP) [1], e.g. when the disease was diagnosed, how long a patient should take some medicines. Due to its importance for many other NLP tasks, temporal information tagging and annotation from clinical texts have become an active research field (Figure 1 for example). Clinical texts including electronic health records (EHRs) consist of medical history, diagnoses, medications, treatment plans, immunisation dates, allergies, radiology images, and laboratory and test results, which contain a lot of useful information that can provide the opportunity to support medical research [2, 3]. These include the study of the side effects of medications (a.k.a adverse drug effects) taken over time, helping doctors make better decisions for patients, and finding the correlation between drug use and its outcomes, etc. [4, 5].
@@ I Introduction
There are two major sub-tasks: 1) Medication Entity Recognition, which involves finding and identifying named drugs mentioned in unstructured text, the first step toward automating the extraction of medications from EMRs. 2) Automatic Temporal Relation Extraction from free text EMRs. We approach these two steps by applying several advanced DL methods including the BiLSTM+CRF, CNN-BiLSTM, and BERT-base-CNN models to train on the available data sets. Additionally, a SparkNLP tool (DateNormalizer) [6] is used to extract and normalise dates from relative date-time phrases.
II Methodological Design
Our designed methodology includes five main steps as shown in Figure 2, which are Data Pre-processing, Word Embedding, Modelling, Time Expression and Event Extraction, and Temporal Relations and Medications Candidates Classification.
II-A Pre-Processing Details
Clinical texts are mixed with different formats. The target of pre-processing is trying to format different data with certain rules and enrich texts with their potential features. Typically, Pre-processing follows 5 steps: Sentence Segmentation, Tokenisation, Parts of Speech (POS, which explains how a word is used in a sentence) Tagging, and Parsing. For our work, GATE [7], cTAKES [8] and MedSpaCy (a new clinical text processing toolkit in Python) shows excellent power in building rule-based modules to retrieve extra features. Additionally, the library ‘spacy’ is valuable and powerful in linking mentions to the associated verb.
II-B Word Embeddings
Word embedding techniques play a role in capturing the meaning, semantic relationship, and context of different words while creating word representations [9]. A word embedding is a technique used to generate a dense vector representation of words that include context words about their own. In addition to one-hot embedding that results in sparse matrix embedding, there are enhanced versions of simple bag-of-words models, such as word counts and frequency counters, that essentially represent sparse vectors.
II-B1 One-hot Encoding
A popular encoding technique is called “one-hot encoding”. In NLP, if a dictionary or sequence has N fields, each field can be represented by an n-dimensional one-hot vector. One-hot encoding can convert classified data into a unified digital format, which is convenient for machine learning algorithms to process and calculate since computers cannot understand non-numeric data. And the fixed dimension vector is convenient for the machine learning algorithm to calculate the linear algebra. Data that don’t relate to one another can benefit from one-hot encoding.
II-B2 Bag-of-words embedding
A bag-of-words model (BoW) is a way of extracting features from the text for use in modelling, such as with machine learning algorithms, which is a popular way of representing documents [10]. Throughout the realm of information retrieval, the BoW model is used in information retrieval on the assumption that a text may be reduced to a collection of words without regard to its actual order within the document, its grammar, or its syntax. Each word’s presence in the document is autonomous and unrelated to the incidence of other terms. In other words, each word at each place in the document is individually selected without regard to the document’s semantics.
II-B3 GloVe Embedding
GloVe stands for “Global Vectors” [11]. Compared to Word2vec, which only keeps local statistic information, the advantage of GloVe is that it captures both global statistics and local statistics of a corpus, and it is an unsupervised learning pre-trained model that represents words in sentences. It has the ability to extract semantic relationships. The main idea of the Glove method is that “semantic relationships between words can be extracted from the co-occurrence matrix” [12]. Given a document having a sentence of n words, the co-occurrence matrix X will be a n*n matrix.
@@ II-B3 GloVe Embedding
II-C Deployed Learning Models
II-C1 BiLSTM
@@
In machine learning fields, the problem that the model cannot get the previous moment information due to the vanishing gradient is also called the long-term dependence problem. In order to solve this problem, many researchers have tried to improve the Recurrent Neural Network (RNN) model, including BRNN [13], GRU [14], and LSTM [15], etc., among which the LSTM model is the most widely used one.
@@
LSTM is an improved RNN structure that can effectively avoid the long-term dependence problem that the traditional RNNs cannot solve. In the later work, many people adjusted and popularised the model, and the modern, widely used LSTM structure came into being. LSTM network has the same chain structure as traditional RNNs, but the repeating module has a different design. The RNN model only has a single layer of repeated modules. In contrast, the LSTM network has four interacting neural network layers, which makes the LSTM model more complex and sacrifices a certain amount of computational time, but the model output becomes better. The standard LSTM model mainly includes three groups of adaptive element multiplication gates, namely, the input gate, forgetting gate, and output gate.
@@
In actual text processing applications, the text is characterised by sequential feature correlation, and words in the current moment are not only affected by words in the past moment but also by words in the future moment. In order to make the model contain contextual information at the same time, researchers put forward a bidirectional LSTM model (BiLSTM) [16, 17]. BiLSTM contains two LSTM layers, which train the forward and backward sequences, respectively, and both LSTM layers are connected to the output layer. This forward-backward two-layer LSTM structure can provide historical information and future information to the output layer at the same time, which can not only retain the advantages of LSTM to solve long-term dependence but also take into account the context information and effectively deal with sequence problems.
II-C2 CRF model
@@
Conditional random fields (CRFs) [18] are a type of statistical modelling methodology frequently employed in pattern recognition and machine learning, as well as for structured prediction, which is a widely-applied modelling strategy for many NLP tasks [19, 20]. Data having an underlying graph structure can be better modelled using CRFs, a graphical model that can take advantage of structural dependencies between outputs. In contrast to classifiers, which make label predictions for isolated data without considering their “neighbouring” samples, CRFs are able to account for context [21, 22, 23]. In order to accomplish this goal, the predictions are modelled in the form of a graphical model, which depicts the existence of relations between the predictions. The choice of the graph to implement is context-specific. In NLP fields, for instance, “linear chain” CRFs are prevalent because each prediction depends only on its immediate neighbours. Connecting neighbouring and/or analogous regions ensures that all regions within a picture are predicted in the same way by the graph.
II-C3 CNN model
@@
In deep learning (DL), a classic convolutional neural network (CNN) is a class of artificial neural networks (ANN), most commonly applied to analyse visual imagery [24]. A CNN model mainly consists of these layers: input layer, convolution layer, relu layer, pooling layer and full connection layer. One of CNN’s most alluring quality is its ability to make use of spatial or temporal correlation in data. CNN is divided into multiple learning phases, each of which is made up of a combination of convolutional layers, nonlinear processing units, and subsampling layers [25]. In a CNN, each layer of the network uses a set of convolutional kernels to perform many transformations. The convolution process simplifies extracting useful features from spatially correlated data points. We will use CNN to learn the character relations within word tokens for our task.
II-C4 BERT model
@@
Brief for “Bidirectional Encoder Representations from Transformers”, BERT [26] is founded on Transformer structure [27], a cutting-edge DL model in which all output elements are related to all input elements and the weights between them are dynamically computed based on their connections. BERT is unique in that it can be used with text that is read either from left to right or right to left. With this characteristic, BERT was originally taught for two NLP tasks: Next Sentence Prediction and Masked Language Modelling. The input for BERT is comprised of Token Embeddings, Segment Embeddings, and Position Embeddings, followed by pre-training. BERT enables effective dynamic feature encoding of polysemous words. In other words, the same word might produce a variety of word vector outputs depending on the language environment. Additionally, the pre-trained model’s embedding layer transfers the knowledge from the corpus to make the word vectors more generic, greatly enhancing the model’s accuracy.
II-D Time Expressions and Events Named Entity Recognition
@@
The goal of NER task is to mark up temporal information present in clinical text in order to enable reasoning on relevant events (medications) for each patient. In this work, a commonly used format in entity tagging “Inside–outside–beginning (IOB)” [28], is used, which is explained below:
@@
II-E Temporal Relations and Medication Candidates Classification
Relation extraction is the task of predicting attributes and relations for entities in a sentence. Extracted relations usually occur between two or more entities of a particular type (e.g. Medication, Dosage) and fall into several semantic categories. In this work, these relations that are discussed are between medications and corresponding dates. For instance, in the example text from Figure 1, there are medication “Methotrexate” that happened between “May 2019” and“February 2020”. Generally, there are eight types of temporal relations, including before, after, simultaneous, overlap, begun_by, ended_by, during, and before_overlap suggested by [29]. However, only the types of before, after, and overlap meet our task. A BERT-based model with CNN as an additional learning layer is implemented for this step.
II-F Post-processing
After the phrases of entity recognition and relation extraction, to make the result more robust and representative, this post-processing step is to convert results to a structured table or CSV shape. However, in the previous 2012 i2b2 NLP challenge on temporal relation extraction, post-processing was considered a set of classification routines for the EVENTs only, and the TIMEx catergory remained unclassified [30]. In this work, we apply an open resource tool SparkNLP [^1] https://github.com/JohnSnowLabs/spark-nlp, combined with our medication entity recognition and temporal relation extraction system to generate more informative results and a structured representation table e.g. the example Table I generated from the example text in Figure 1.
TABLE I: Extraction Result Example
@@ II-F Post-processing
III Experimental Evaluations
III-A Dataset and Evaluation Metrics
In this work, the i2b2 challenge track 2009 and 2012 datasets were used. The objective is to gather details about every drug that the patient is known to take for each patient report that is given. Partners Healthcare’s discharge summaries will be the input for the medication challenge [29]. These files are loosely organised. Many drugs are covered in the narrative text in addition to being listed by name in lists of pharmaceuticals. According to statistical analysis, it contains 447 EMRs, with 252 annotated, which could be used to train and test the model. While the data from the i2b2-2012 temporal relations challenge has similar content as the challenge track of 2009 does, it has a different format, additional entity categories, and temporal relations between events. It consists of 760 files including 190 TXT files (texts), 190 XML.EXTENT files (annotation and position in original text), XML files (annotation and text) and TLINK files (relations) [31]. The data sets are divided into 70%, 15% and 15% for training, validation, and testing, respectively on both these two i2b2 challenges. The entity tags used in the i2b2 challenge track 2009 and 2012 have some differences which are shown in Table II and Table III.
TABLE II: Tags Used in the i2b2 2009 Dataset.
Tag Meaning Example
m medication Percocet
do dosage 3.2mg
f frequency twice a day
mo mode (/route of administration) Mode: “nm” (not “Tablet”)
du duration 10-day course
r reasons Dizziness
TABLE III: Tags Used in the i2b2 2012 Dataset.
Tag Meaning Example
CLINICAL_DEPT clinical department emergency room
EVIDENTIAL events that have an ‘evidential’ nature CT shows
TEST clinical tests CT
PROBLEM symptoms sickness
TREATMENT medications, surgeries and other procedures Levaquin
The evaluation metrics we used are Precision, Recall, F1 score, and Accuracy based on exact matching. Considering that there are multiple categories of labels, two different kinds of evaluation criteria are applied: 1) macro average evaluation, the function to compute precision, recall and F1-score, for each label and returns the average without considering the proportion for each label in the dataset; and 2) weighted average evaluation, the function to compute precision, recall and F1-score for each label, and returns the average considering the proportion for each label in the dataset.
III-B Model-I: BiLSTM-CRF for NER Task
Combined with the DL model BiLSTM and graph model CRF, we carried out the NER task of electronic diseases using BiLSTM-CRF. The specific steps carried out are as follows:
@@ III-B Model-I: BiLSTM-CRF for NER Task
Using one-hot word embedding, the evaluation scores on 2009 and 2012 data sets are shown in Table V and VI with the explanations of each tag in Table II and III. The model parameters for 2012 data is listed in Table IV. The model learning parameters for 2009 data is the same except for Max Epochs which is set to be 10, and the number of total trainable parameters is 1,292,188 instead of 506,138. The learning curves on Accuracy and Loss using 2009 and 2012 data are displayed in Figure 3 and 4, which have similar patterns.
TABLE IV: Model Parameters of BiLSTM+CRF for the 2012 Dataset
parameter value
DENSE_EMBEDDING 50
LSTM_UNITS 50
LSTM_DROPOUT 0.2
DENSE_UNITS 100
BATCH_SIZE 256
MAX_EPOCHS 30
LEARNING_RATE 0.0001
Total Param 506,138
Figure 3: BiLSTM-CRF Model Training Process on i2b2-2009 Dataset
Figure 4: BiLSTM-CRF Model Training Process on i2b2-2012 Dataset
TABLE V: NER Evaluation on the 2009 Dataset with BiLSTM-CRF.
@@
TABLE VI: NER Evaluation on the 2012 Dataset with BiLSTM-CRF.
@@
As mentioned, evaluation metrics were based on exact-matching criteria, and with this setting, the BiLSTM+CRF model achieves a 98.50 weighted average F1 and 58.55 of Macro average F1 on 2009 data. Apparently, combined with the evaluation scores of each subcategory, it can be seen that the amount of entities of each category varies significantly due to the uneven data distribution, which leads to the scores between entity categories varying wildly.
As demonstrated in Table V, the BiLSTM+CRF model has better prediction performances on class “B-mo”, “B-f”, “I-do”, “I-m”, “B-do” in addition to “o” with 0.75+ F1. and “B-m” in the 2009 dataset. The classes that have under 0.50 F1 scores also have much fewer supporting labels, less than 200. Similar patterns can be observed in the experiments on 2012 data set, i.e. higher number of supporting labels produce high F1 scores.
However, there are some kind of exceptions, for instance, “I-r” and “B-du” have the supporting labels (141, 72) but their corresponding Macro Avg scores are (16.95, 24.44). In addition, the results based on precision, recall, and F1-score are very different. For most sub-classes, the model has higher precision than recall values which means it returns fewer results, but most of its predicted labels are correct when compared to the gold labels. Namely, the percentage of Negative labels is lower.
In order to explore the impacts of different embedding inputs for the BiLSTM+CRF model, we trained the model with Glove (glove.6B.50d), word2vector, in addition to one-hot encoding, and the average and standard deviations of metrics are recorded. Table VII demonstrates the performance of each model on the same dataset using Macro average F1 score. Glove embedding achieved 72.48% of precision, 52.58% of recall, and 60.86% of f1-score, respectively which increased by around 2% than one-hot embedding. For word2vec embedding, this model achieved exact 73.62% precision, 60.34% of recall, and 61.48% of f1-score achieving a further improvement of 1.14% in precision, 7.76% in recall and 0.62% than Glove embedding.
TABLE VII: Macro Avg Results of Different Embeddings on the 2009 Test Set using BiLSTM+CRF.
@@
TABLE VIII: Macro Avg Results of Experiments on the 2009 Test Dataset with CNN-BiLSTM.
@@
III-C Model-II: CNN-BiLSTM for NER